# Efficient Test Data Generation for MC/DC with OCL and Search

### v1

@@ 5.3 Threats to Validity
Like any empirical evaluation, our experiment is also subjected to different threats that can affect its validity. The possible threats to the validity of our experiment are discussed individually in the following subsections.
@@ 5.3.3 Conclusion Validity Threat
The threat to conclusion validity occurs when treatments can impact the outcome of the experimental results. To minimize the effect of this threat, we repeated all OCL constraints 100 times for each treatment. We analyzed the statistical significance of experimental results based on the success rates using Fisher’s exact test and iteration counts using the Wilcoxon test. We also measured the effect size by calculating Odds ratios and using Vargha-Delaney’s Â${}_{12}$ measure [99]. We analyzed the outcomes of the comparison with existing constraint solvers based on the percentage of successfully solved constraints and the average time.
@@ 6.1 Code-based Approaches
In the work related to code-based approaches, some approaches target branch coverage [47, 48], path coverage [63], and multiple conditions coverage [39] for generating test data. A few approaches target MC/DC criterion such as  [110, 104, 11]. Godboley et al. [40] proposed an approach to improve MC/DC coverage using code transformations. Das and Mall [31] proposed a technique that transforms code under test to achieve MC/DC coverage. Li et al. [66] proposed an approach to improve MC/DC coverage and fault detection effectiveness using combinatorial testing. In addition to the previously mentioned approaches, some approaches aim at addressing multiple coverage criteria and incorporate the concept of solutions archiving. Rojas et al. [85] performed an empirical study to analyze the coverage of the program under test using the whole test suite generation approach with archiving or covered targets. Similarly, Panichella et al. [77] proposed a many-objective search algorithm (DynaMOSA) that dynamically selects coverage targets for test case generation. Cegin and Rastocny [22] proposed an approach that uses reinforcement learning to maximize the MC/DC coverage of the code under test. In recent work, Godboley et al. [41] presented a method to optimize input programs’ MC/DC coverage of input programs. The main advantage of our methods over the above-mentioned approaches is that our approach uses UML models and OCL constraints to generate test data and is independent of source code written in any programming language. Furthermore, OCL and UML models are used to generate source code [24, 70], in which OCL constraints are translated into conditions in code [54]. Our strategy can be applied to achieve MC/DC coverage of the source code generated from these models. While this approach can achieve sufficient code coverage, it is important to note that similar coverage cannot be guaranteed, as reported by Hemmati et al. [54].
@@ 6.2 Specification-based Approaches
In a recent work [95], a hybrid test data generation approach is proposed that utilizes the search and SMT for solving OCL constraints. The key difference is that our work supports the MC/DC coverage of OCL constraints while generating test data. Hemmati et al. [54] evaluated the use of MC/DC criterion on guard conditions of the state machine for the generation of coverage-based test data. The evaluation results indicated that the search budget, search space, and execution time are the main limitations of solving industrial constraints with search techniques. Motivated by these findings, we introduced an MC/DC test data generation strategy featuring CBR and range reduction methods designed to optimize the search process.
@@
SAT/Alloy. Przigoda et al. [79, 80, 81] introduced verification of structural and behavioral UML/OCL models in symbolic form using SAT solvers. Soeken et al. [94] demonstrated the encoding of OCL constraints into bit-vector formats and the verification of UML/OCL models using SAT solvers. Wotawa et al. [105] proposed an approach that transforms code into constraints, test case generation into a constraint satisfaction problem, and use MINION [38] to solve generate distinct tests. Van et al. [98] introduced a UML-CASTING tool that generates test cases by using UML models and solving OCL constraints. Krieger et al. [62] presented an approach that converts operation contracts into Boolean formulas and solves those constraints using a SAT solver.
@@
CSP. Ulke et al. [96] proposed a method for partial evaluation of OCL constraints by formulating them as CSP problems. Gogolla et al. [43, 44] studied UML/OCL models verification and validation with the USE tool. Gogolla and Stüber [42] developed a metric for OCL constraints verification based on expert study. Ferdjoukh et al. [34] presented a model generation approach that converts metamodel into CSP and uses CPS solver to check conformance. Gonzalez et al. [45] introduced an EMFtoCSP tool that provides automatic verification of EMF models by transforming them into CPS problems. Hilken et al. [55] proposed an approach for transforming UML/OCL models into based models and using CSP solvers for verification. Aichernig et al. [3] proposed an approach that transforms test cases into CSP and performs OCL mutations to generate fault-oriented test cases. Cabot et al. [18, 19] proposed an approach for UML/OCL models verification using CSP solvers. Packevicius et al. [76] presented an approach for generating test data targeting complex data types in model constraints. Erata et al. [33] introduced AlloyInEcore which identifies inconsistent models and fills in partial models based on metamodel semantics. Bordbar and Anastasakis [15] presented the UML2Alloy tool that converts the UML model into Alloy and uses a CSP solver to verify OCL constraints.
@@
In all the aforementioned approaches related to CSP, SAT/Alloy, SMT, and theorem provers, our work distinguishes itself in three main aspects. First, our work addresses the lack of support for OCL collection operations by these works, which are common in industrial-level constraints. Second, unlike these approaches that require translating OCL constraints into intermediate formalisms (e.g., DNF), our method avoids this additional step to facilitate industrial application. Third, none of the mentioned approaches support test data generation with a focus on achieving MC/DC coverage of OCL constraints, whereas our approach specifically targets this criterion.
@@ 6.4 Domain Reduction Techniques
Harman et al. [53] proposed a program slicing-based input domain reduction technique to generate test data for aspect-oriented programs. Their approach operates at the source code level and is specific to aspect-oriented programs, while our methods are designed for handling OCL constraints. Harman et al. [51] presented a theoretical and empirical analysis to investigate the impact of search space reduction on test data generation. The results of both types of analysis indicate that reducing search space enhances the performance of search algorithms. Arcuri and Yao [10] introduced search space reduction for Java containers to minimize the computation cost of search algorithms. Ribeiro et al. [83] introduced the input domain reduction strategy to generate tests for object-oriented Java programs. Rojas et al. [84] analyzed different seeding strategies that are used to generate test cases for Java programs. The works mentioned above are specifically designed for Java programs, while our methods are focused on reducing the range for MC/DC constraint solving. Clariso et al. [26] presented a technique that focuses on assisting users in selecting bounds for UML/OCL models verification, whereas our range reduction method automatically computes ranges for MC/DC constraints to facilitate test data generation. Bidgoli and Haghighi [14] proposed an approach to reduce search space for a program input variable. Their approach is confined to individual clauses, whereas our range reduction method considers all independent and dependent clauses used in MC/DC constraints, which is essential for generating test data that adheres to MC/DC combinations. Das and Pratihar [30] proposed a dynamic search space reduction approach for genetic algorithms based on the diversity of the population. In comparison, our range reduction method computes ranges for MC/DC constraints at the initial stage and applies these reduced ranges before starting the search process. This flexibility enables utilizing our range reduction method with any search algorithm.

### v2

@@ 5.3 Threats to Validity
Like any empirical evaluation, our experiment is also subjected to different threats that can affect its validity. The possible threats to the validity of our experiment and their mitigation approaches are discussed individually in the following subsections.
@@ 5.3.3 Conclusion Validity Threat
The threat to conclusion validity occurs when treatments can impact the outcome of the experimental results. To minimize the effect of this threat, we repeated all OCL constraints 100 times for each treatment. We analyzed the statistical significance of experimental results based on the success rates using Fisher’s exact test and iteration counts using the Wilcoxon test. We also measured the effect size by calculating Odds ratios and using Vargha-Delaney’s Â12 measure [99]. We analyzed the outcomes of the comparison with existing constraint solvers based on the percentage of successfully solved constraints and the average time.
@@ 6.1 Code-based Approaches
In works related to code-based approaches, some approaches target branch coverage [47, 48], path coverage [63], and multiple conditions coverage [39] for generating test data. A few approaches also target MC/DC criterion  [110, 104, 11]. In this regard, Godboley et al. [40] proposed an approach to improve MC/DC coverage using code transformations. Das and Mall [31] proposed a technique that transforms code under test to achieve MC/DC coverage. Li et al. [66] proposed an approach to improve MC/DC coverage and fault detection effectiveness using combinatorial testing. In addition to the previously mentioned approaches, some approaches aim at addressing multiple coverage criteria and incorporate the solutions archiving concept. Rojas et al. [85] performed an empirical study to analyze the coverage of the program under test using the whole test suite generation approach with archiving or covered targets. Similarly, Panichella et al. [77] proposed a many-objective search algorithm (DynaMOSA) that dynamically selects coverage targets for test case generation. Cegin and Rastocny [22] proposed an approach that uses reinforcement learning to maximize the MC/DC coverage of the code under test. In recent work, Godboley et al. [41] presented a method to optimize input programs’ MC/DC coverage of input programs. The main advantage of our methods over the above-mentioned approaches is that our approach uses UML models and OCL constraints to generate test data and is independent of source code written in any programming language. Furthermore, OCL and UML models are used to generate source code [24, 70], in which OCL constraints are translated into conditions in code [54]. Our strategy can be applied to achieve MC/DC coverage of the source code generated from these models. While this approach can achieve sufficient code coverage, it is important to note that similar coverage cannot be guaranteed, as reported by Hemmati et al. [54].
@@ 6.2 Specification-based Approaches
In a recent work [95], a hybrid test data generation approach (PLEDGE) is proposed that utilizes the search and SMT for solving OCL constraints. The key difference is that our work supports the MC/DC coverage of OCL constraints while generating test data. Hemmati et al. [54] evaluated the use of MC/DC criterion on guard conditions of the state machine for the generation of coverage-based test data. The evaluation results indicated that the search budget, search space, and execution time are the main limitations of solving industrial constraints with search techniques. Motivated by these findings, we introduced an MC/DC test data generation strategy featuring CBR and range reduction methods designed to optimize the search process.
@@
SAT/Alloy. Przigoda et al. [79, 80, 81] introduced verification of structural and behavioral UML/OCL models in symbolic form using SAT solvers. Soeken et al. [94] demonstrated the encoding of OCL constraints into bit-vector formats and the verification of UML/OCL models using SAT solvers. Wotawa et al. [105] proposed an approach that transforms code into constraints, test case generation into a constraint satisfaction problem, and uses MINION [38] to solve generate distinct tests. Van et al. [98] introduced a UML-CASTING tool that generates test cases by using UML models and solving OCL constraints. Krieger et al. [62] presented an approach that converts operation contracts into Boolean formulas and solves those constraints using the SAT solver.
@@
CSP. Ulke et al. [96] proposed a method for partial evaluation of OCL constraints by formulating them as CSP problems. Gogolla et al. [43, 44] studied UML/OCL models verification and validation with the USE tool. Gogolla and Stüber [42] developed a metric for OCL constraints verification based on an expert study. Ferdjoukh et al. [34] presented a model generation approach that converts metamodel into CSP and uses a CPS solver to check conformance. Gonzalez et al. [45] introduced an EMFtoCSP tool that provides automatic verification of EMF models by transforming them into CPS problems. Hilken et al. [55] proposed an approach for transforming UML/OCL models into based models and using CSP solvers for verification. Aichernig et al. [3] proposed an approach that transforms test cases into CSP and performs OCL mutations to generate fault-oriented test cases. Cabot et al. [18, 19] proposed an approach (UMLtoCSP) for UML/OCL model verification using CSP solvers. Packevicius et al. [76] presented an approach for generating test data targeting complex data types in model constraints. Erata et al. [33] introduced AlloyInEcore which identifies inconsistent models and fills in partial models based on metamodel semantics. Bordbar and Anastasakis [15] presented the UML2Alloy tool that converts the UML model into Alloy and uses a CSP solver to verify OCL constraints.
@@
In all the aforementioned approaches related to CSP, SAT/Alloy, SMT, and theorem provers, our work distinguishes itself in two main aspects. First, our work addresses the lack of support for OCL collection operations by these works, which are common in industrial-level constraints. Second, none of the mentioned approaches support test data generation with a focus on achieving MC/DC coverage of OCL constraints, whereas our approach specifically targets this criterion.
@@ 6.4 Domain Reduction Techniques
Harman et al. [53] proposed a program slicing-based input domain reduction technique to generate test data for aspect-oriented programs. Their approach operates at the source code level and is specific to aspect-oriented programs, while our methods are designed for handling OCL constraints. Harman et al. [51] presented a theoretical and empirical analysis to investigate the impact of search space reduction on test data generation. The results of both types of analysis indicate that reducing search space enhances the performance of search algorithms. Arcuri and Yao [10] introduced search space reduction for Java containers to minimize the computation cost of search algorithms. Ribeiro et al. [83] introduced the input domain reduction strategy to generate tests for object-oriented Java programs. Rojas et al. [84] analyzed different seeding strategies that are used to generate test cases for Java programs. The works mentioned above are specifically designed for Java programs, while our methods are focused on reducing the range for MC/DC constraint solving. Clariso et al. [26] presented a technique that focuses on assisting users in selecting bounds for UML/OCL model verification, whereas our range reduction method automatically computes ranges for MC/DC constraints to facilitate test data generation. Bidgoli and Haghighi [14] proposed an approach to reduce search space for a program input variable. Their approach is confined to individual clauses, whereas our range reduction method considers all independent and dependent clauses used in MC/DC constraints, which is essential for generating test data that adheres to MC/DC combinations. Das and Pratihar [30] proposed a dynamic search space reduction approach for genetic algorithms based on the diversity of the population. In comparison, our range reduction method computes ranges for MC/DC constraints at the initial stage and applies these reduced ranges before starting the search process. This flexibility enables utilizing our range reduction method with any search algorithm.